Shift report 20010830 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Aug 30 2001 - 23:42:29 EDT

  • Next message: Apache: "Shift report 20010831 00:00-08:00"

    ------------------------------------------------------------
    Date: Thu Aug 30 23:42:29 2001, Shift: 16:00-24:00
    Supervisor: R. Debbe, with: None whatsoever!!
    Log book#8 pages 44 - 46
    
    Shift Summary:
    ------------------------------------------------------------
    Machine status: Great store started at 15:30 with 14 10^9 in blue and 22 10^9 in yellow
    the same store is running now.
    Achievements: Data runs 4504,4506, 4507 and 4508 (still running)
    I have no way to estimate the total number of trigger 1 in this setting.
    Whithin the runs acquired in this shift there are: 500 K
    Dana was in when I arrived, he was trying to match gains in the ZDC
    but left without changing anything.
    
    Detector problems: Not much two TDC channels missing in H1 15 and 33
    I also found the flow of argon in T1 and T2 was way too low, details in log book pg 45
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Aug 30 2001 - 23:42:59 EDT